OpenStudy (anonymous):

Rationalize the denominator. The answer should be expressed in simplified form.

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ \sqrt{11}-\sqrt{5} }\]

OpenStudy (anonymous):

\[\frac{ \sqrt{10} }{ 6 }+\frac{ \sqrt{22} }{ 6 }\] Is that correct?

OpenStudy (anonymous):

@iambatman

OpenStudy (solomonzelman):

for question 1, multiply by a conjugate top and bottom. Also, iambatman is very very smart, but he is offline right now.

OpenStudy (solomonzelman):

In this conjugate is \(\large\color{black}{ \bf \sqrt{11}+\sqrt{5} }\)

OpenStudy (anonymous):

How did it become plus?

OpenStudy (solomonzelman):

it didn't \(\large\color{red}{ \bf become }\) plus, I am using a conjugate to rationalize the denominator.

OpenStudy (anonymous):

Oooh, I get it now

OpenStudy (solomonzelman):

SO tell me what do you get when you multiply top and bottom times \(\large\color{brown }{ \bf \sqrt{11}+\sqrt{5} }\) ?

OpenStudy (anonymous):

\[\sqrt{22} + \sqrt{10}\] The bottom would just be: 11 & 5

OpenStudy (solomonzelman):

are actually correct, I thought it was another problem... duh, I am so dumb ! And the bottom would be 11-5, b/c \(\large\color{black}{ \bf (a-b)(a+b)=a^2 }\) \(\large\color{red}{ \bf - }\) \(\large\color{black}{ \bf b^2 }\)

OpenStudy (solomonzelman):

and 11-5 is 6.

OpenStudy (anonymous):

So, my answer is correct?

OpenStudy (anonymous):

You are NOT dumb, at all lol

OpenStudy (solomonzelman):

Yes, your answer is correct :)

OpenStudy (anonymous):

O r can i just put both of the numerators together on top & just have one 6 on the bottom or does it have to be like that?

OpenStudy (solomonzelman):

\(\large\color{black}{ \bf \frac{\Huge \sqrt{10}+\sqrt{22} }{\Huge6} }\) and what you have is the same exact thing, but I think it is a tiny bit better to write them as one fraction.

OpenStudy (anonymous):

Ok, thank you!